OpenStudy (anonymous):

∫pi/2 between 0 of : e^(cos x) sin x dx ? I GIVE MEDALS

OpenStudy (anonymous):

let u = cosx du = -sinx dx

OpenStudy (anonymous):

yeah I already did this :) but after ?

OpenStudy (ikram002p):

after that integrate ! integral -e^u du

OpenStudy (anonymous):

yeah but it doesnt work :(

OpenStudy (anonymous):

cuz the answer is :e-1= 1,72

OpenStudy (anonymous):

but cos of pi/2 = 0 .... :(

OpenStudy (ikram002p):

this is ur integral right ? \[\int\limits_{0}^{\frac{ \pi }{2 }} e^\cosx \sin x dx \]

OpenStudy (anonymous):

yes!

OpenStudy (ikram002p):

let u = cosx du = -sinx dx as sourwing said

OpenStudy (anonymous):

yes Thats what I did

OpenStudy (ikram002p):

\[\int\limits_{ }^{ }-e^u du =-e^u =- e^\cosx \]

OpenStudy (ikram002p):

from 0 to pi/2

OpenStudy (anonymous):

yeah

OpenStudy (ikram002p):

\[-e^cospi/2 +e^ \cos0\]

OpenStudy (anonymous):

- isn't for all the equation?

OpenStudy (anonymous):

ah oops sorry yeah

OpenStudy (ikram002p):

-e^0 +e^1=e-1

OpenStudy (ikram002p):

e= ?

OpenStudy (anonymous):

omg I see now

OpenStudy (ikram002p):

^^ im a bit sleepy gn :D

OpenStudy (anonymous):

My god , now, I dont know how I could thinks it was difficult hahaha :) good nught :)!!

OpenStudy (ikram002p):

na its not diffecult lol by the way are u arabian ?

OpenStudy (anonymous):

hahaha nooo! Canadian :)

OpenStudy (anonymous):

But french :s

OpenStudy (ikram002p):

lol sry i told u im a bit sleepy XD its 2:12 am So Good night ^^

OpenStudy (anonymous):

you yes?:o its 7h13 here !!